OpenStudy (anonymous):

A IRS auditor randomly selects 3 tax returns from 60 returns of which 11 contain errors. What is the probability that she selects none of those containing errors? Round to four decimal places.

OpenStudy (kropot72):

The probability of selecting a return with an error is 11/60. Therefore the probability of selecting a return without an error is: \[\large 1-\frac{11}{60}=\frac{49}{60}\] The probability of selecting 3 tax returns randomly each without an error is given by: \[\large P(3\ without\ errors)=\frac{49}{60}\times\frac{49}{60}\times\frac{49}{60}=you\ can\ calculate\]